# One-Shot Broadcast Joint Source-Channel Coding with Codebook Diversity

### v1

@@ II Problem Setup
We consider a one-shot setting with one encoder and $K$ decoders, where each decoder is connected to the encoder through a private and independent communication channel which maps an input alphabet $\mathcal{X}$ to an output alphabet $\mathcal{Y}$. Furthermore, we restrict our attention to the case where all the channels are identically modeled by a conditional distribution $P_{Y\mid X}$. Under this assumption, the encoder, defined by the mapping $f:\mathcal{W}\to\mathcal{X}$, observes the source symbol $W\in\mathcal{W}$ from $P_{W}$ and broadcasts a message $X\in\mathcal{X}$ over each channel. Subsequently, decoder $k$ views the sample $Y_{k}\in\mathcal{Y}$ at the output of its channel and generates the reconstruction $\hat{Z}_{k}\in\mathcal{Z}$ using a mapping $g:\mathcal{Y}\to\mathcal{Z}$. The probability of excess distortion for an individual decoder under the distortion measure $d:\mathcal{W}\times\mathcal{Z}\to\mathbb{R}_{\geq 0}$ is $P_{ek}=\operatorname{Pr}[d(W,\hat{Z}_{k})>D]$. As we require at least one decoder to recover the source within the specified distortion, the ensemble error probability is then $P_{e}=\operatorname{Pr}[\cap_{i=1}^{K}\{d(W,\hat{Z}_{k})>D\}]$.
We further consider a setting where each decoder has access to a conditionally independent side information realization $T_{k}$ from $P_{T\mid W}$, as motivated in section I. Note that for simplicity we only consider the case where the side information samples seen at the decoders are identically distributed. As opposed to the basic setting, decoder $k$ now generates $\hat{Z}_{k}$ using a mapping $g:\mathcal{Y}\times\mathcal{T}\to\mathcal{Z}$ that can exploit the side information as well as the channel observation. This constitutes a variant of the Wyner-Ziv problem [12] and consequently requires a somewhat different analysis. A block diagram of the full system model with side information is illustrated in fig. 1.
In what follows, we assume that all alphabets, e.g. $\mathcal{W}$, $\mathcal{X}$ and $\mathcal{Y}$, are standard Borel spaces endowed with their respective Borel $\sigma$-algebras and assume a common $\sigma$-finite reference measure $\theta$ on each space. For a pair of measures $\mu$, $\nu$ such that $\nu$ is absolutely continuous with respect to $\mu$, denoted as $\nu\ll\mu$, the Radon-Nikodym derivative is $d\nu/d\mu:\mathcal{U}\to\mathbb{R}_{\geq 0}$. To simplify the exposition, we assume that all probability measures to be discussed are absolutely continuous with respect to the reference measure $\theta$ on the appropriate space and work in terms of the density function $p(u)=(dP/d\theta)(u)$ for any measure $P$. Furthermore, we write $\lambda$ for the Lebesgue measure on $\mathbb{R}$ and $\lambda_{S}$ for its restriction to some $S\subseteq\mathbb{R}$. Logarithms are base-$2$ throughout; the information density for a pair of random variables $(X,Y)\sim P_{X,Y}$ is defined as
@@
[width=]diagrams/system
Figure 1: System model for broadcast JSCC with side information. The source $W$ is encoded to $X$. Decoder $k$ uses the channel output $Y_{k}$ and side information $T_{k}$ to produce its reconstruction $\hat{Z}^{(k)}$.
@@ III-E Hybrid Strategy
$P_{e}\leq\operatorname{\mathbb{E}}\Bigl[\Bigl(1+JP_{Z}(\mathcal{B}_{D}(W)\max_{1\leq k\leq L}2^{\iota_{X;Y}(X;Y_{k})}\Bigr)^{-1}\Bigr].$ (11)
@@ IV List Poisson Matching Lemma
Let $\Pi=\{U_{i},T_{i}\}_{i\in\mathbb{N}}$ be a marked Poisson point process on $\mathcal{U}\times\mathbb{R}_{\geq 0}$ with intensity measure $\mu\times\lambda_{\mathbb{R}_{\geq 0}}$. Further associate with each point of the process an independent discrete random label $Z_{i}$ uniformly distributed on $\{1,\ldots,K\}$ and let $P$, $Q\ll\mu$ be two probability measures on $\mathcal{U}$. Select the indices
@@
and set $U_{p}=U_{K_{p}}$ and $Z_{p}=Z_{K_{p}}$. Given $U_{p}$, let $Y_{1:K}\sim P_{Y\mid X,U}(\cdot\mid X,U_{p})$ iid. We have $(X,U_{p}.Y_{k})\sim P_{X,U,Y}$ for each $k$ and note that the Markov chain condition $Y_{1:K}\leftrightarrow(X,U_{p})\leftrightarrow\{U_{i},T_{i}\}_{i\in\mathbb{N}}$ holds by construction. Next, select
@@
$\leq$ $1-\sum_{k=1}^{K}\biggl(K-\frac{p_{U\mid X}(U_{p}\mid X)}{q_{U\mid Y}(U_{p}\mid Y_{k})}\biggr)^{-1}.$ (17)
@@ V Proof of Theorem 1
First, we generate the full codebook independently from the source $W$ as a marked Poisson point process $\Pi=\{(X_{i},Z_{i}),\tau_{i}\}_{i\in\mathbb{N}}$ on $\mathcal{X}\times\mathcal{Z}\times\mathbb{R}_{\geq 0}$ with intensity measure $P_{X}\times P_{Z}\times\lambda_{\mathbb{R}\geq 0}$, and associate with each point in the codebook a uniform label $C_{i}\in\{1,\ldots K\}$ independent of $\Pi$. Following the construction in [8, Th. 4], we define a modified conditional distribution $P_{\tilde{Z}\mid W}$ to take into account the intersection with the distortion ball $\mathcal{B}_{D}(w)$. The conditional density function is
@@
$\negmedspace{}\biggl(1-\sum_{k=1}^{K}\biggl(K+\frac{p_{X}(X)p_{\tilde{Z}\mid W}(\tilde{Z}\mid W)}{p_{X\mid Y}(X\mid Y_{k})p_{Z}(\tilde{Z})}\biggr)^{-1}\biggr)\biggr]$
@@
$\negmedspace{}\biggl(1-\sum_{k=1}^{K}\bigl(K+p(W)^{-1}2^{-\iota_{X;Y}(X;Y_{k})}\bigr)^{-1}\biggr)\biggr]$
@@
where $(W,X,Y)\sim P_{W}\times P_{X}P_{Y\mid X}$, (a) is by lemma 2 and (b) uses the fact that the $Y_{k}$’s are identically distributed given $X$ and the identity $\operatorname{\mathbb{E}}[\operatorname{\mathds{1}}\{\rho(W)>0\}]=\operatorname{Pr}[\rho>0]$. ∎
@@ VI Proof of Theorem 2
We now establish the bound with side information in theorem 2. Since the side information is not visible to the encoder, it cannot restrict the codeword to the permissible distortion ball as in the proof of theorem 1. Let the codebook be a marked Poisson point process $\Pi=\{(X_{i},U_{i}),\tau_{i}\}_{i\in\mathbb{N}}$ on $\mathcal{X}\times\mathcal{U}\times\mathbb{R}_{\geq 0}$ with intensity measure $P_{X}\times P_{U}\times\lambda_{\mathbb{R}\geq 0}$ independent of $W$, where $P_{U}$ is the marginal of $P_{W}P_{U\mid W}$. We associate an independent uniform label $C_{i}\in\{1,\ldots K\}$ with each point. The encoder observes $W$ and selects
@@ VIII Conclusion
We derived one-shot achievability results for a broadcast JSCC problem in a distributed setting where success is determined by most favorable outcome among $K$ decoders. We identified two competing ways in which having multiple decoders can reduce the error probability, namely channel diversity, achieved by sharing a codebook and taking multiple independent looks at the channel, and codebook diversity, which requires forcing the decoders’ codebooks to be disjoint. During our analysis, we developed the list PML as an extension of the PML in order to obtain bounds on the error probability. Moreover, we derived a hybrid coding scheme that optimally balances channel and codebook diversity; numerical results demonstrated that it improves on approaches that rely on one form of diversity alone.
@@ Appendix A Proof of Lemma 1
Thus, an upper bound for $\operatorname{Pr}[U_{p}\neq U_{q1}\mid U_{p}=u,Z_{p}=1]$ is the probability of $\Pi^{\prime}$ seeing at least one arrival in the region
@@
$\leq$ $\frac{t}{K}\int_{\mathcal{U}}\frac{p(u)q(v)}{q(u)}\mu(dv)=\frac{tp(u)}{q(u)}.$ (48)
@@
$=$ $\operatorname{Pr}[N>0\mid U_{p}=u,Z_{p}=1]\leq 1-\biggl(1+\frac{p(u)}{Kq(u)}\biggr)^{-1}$ (52)
@@ Appendix B Proof of Lemma 2
and set $U_{p}=U_{K_{p}}$, $Z_{p}=Z_{K_{p}}$. Given $U_{p}$, let $Y_{1:K}\sim P_{Y\mid X,U}(\cdot\mid X,U_{p})$ iid. We have $(X,U_{p}.Y_{k})\sim P_{X,U,Y}$ for each $k$ and note that the Markov chain condition $Y_{1:K}\leftrightarrow(X,U_{p})\leftrightarrow\{U_{i},T_{i}\}_{i\in\mathbb{N}}$ holds by construction. Next, select
@@
In what follows, condition on $\tilde{T}_{p}=t$, $U_{p}=u$, $Z_{p}=1$ and $Y_{1:K}=y_{1:K}$. Mirroring the arguments used to prove lemma 1, $\tilde{T}_{i}\geq\tilde{T}_{p}=t$ for each $i\neq K_{p}$ and therefore, by excluding the first arrival at time $t$, $\Pi^{\prime}=\{U_{i},\tilde{T}_{i}\}_{i\neq K_{p}}$ becomes a marked Poisson point process with intensity measure $P_{U\mid X}(\cdot\mid X)\times\lambda_{[t,\infty)}$ independent of $U_{p}$ and each $Z_{i}$. By the thinning property, $\Pi_{1}^{\prime}=\{U_{i},\tilde{T}_{i}\}_{i\neq K_{p},Z_{i}=1}$ is then a marked Poisson point process with intensity measure $(P_{U\mid X}(\cdot\mid X)\times\lambda_{[t,\infty)})/J$. By the Markov chain condition noted earlier, both $\Pi^{\prime}$ and $\Pi_{1}^{\prime}$ are also conditionally independent of $Y_{1:K}$ given $X$ and $U_{p}$. As in the unconditional case, assume $p_{U\mid X}(u\mid X)>0$. First consider that $q_{U\mid Y}(u\mid Y_{k})>0$ for at least one $k\in\pi(1)$. Selecting the same indices implies the selected values are the same, and consequently
@@
$\mathcal{B}=\left\{(v,s):s<\min_{1\leq k\leq}\frac{tp_{U\mid X}(u\mid X)q_{U\mid Y}(v\mid Y_{k})}{q_{U\mid Y}(u\mid Y_{k})p_{U\mid X}(v\mid X)}\right\}.$ (64)
@@
$=$ $\frac{t}{J}\int_{\mathcal{U}}\min_{1\leq k\leq L}\max\left\{0,\frac{p_{U\mid X}(u\mid X)}{q_{U\mid Y}(u\mid Y_{k})}-\frac{p_{U\mid X}(v\mid X)}{q_{U\mid Y}(u\mid Y_{k})}\right\}$
@@
$\geq$ $\exp\left(-\frac{t}{J}\min_{1\leq k\leq K}\frac{p_{U\mid X}(u\mid X)}{q_{U\mid Y}(u\mid Y_{k})}\right).$ (73)
@@ Appendix C Error Probability of the Hybrid Scheme
In this section, we provide a proof of (11) via lemma 3. By setting $J=1$, this also serves to show the bound in (9) for the baseline scheme’s performance. Following the approach used in the proof of theorem 1, let the codebook be a marked Poisson point process $\Pi=\{(X_{i},Z_{i}),\tau_{i}\}_{i\in\mathbb{N}}$ on $\mathcal{X}\times\mathcal{Z}\times\mathbb{R}_{\geq 0}$ with intensity measure $P_{X}\times P_{Z}\times\lambda_{\mathbb{R}\geq 0}$ independent of the source $W$. Associate with each point a label $C_{i}$ uniform on $\{1,\ldots,J\}$ independent of $\Pi$. The encoder uses the modified conditional distribution $P_{\tilde{Z}\mid W}$ as in (18) and selects its index $K_{p}$ using (19). Meanwhile, decoder $k$ selects $K_{qk}$ using (10). As before, let $X=X_{K_{p}}$, $\tilde{Z}=Z_{K_{p}}$, $\hat{Z}_{k}=Z_{K_{qk}}$ and $\hat{X}_{k}=\hat{X}_{K_{qk}}$. The channel outputs observed by the decoders follow $Y_{1:K}\sim P_{Y\mid X}(\cdot\mid X)$ iid. We have $(W,\tilde{Z},X,Y_{k})\sim P_{W}P_{\tilde{Z}\mid W}\times P_{X}P_{Y\mid X}$ for each $k$ by lemma 3. Additionally, define $\pi(j)=\{(j-1)L+1,\ldots,jL\}$ for $1\leq j\leq J$. Then,
@@
$\negmedspace{}\times\biggl(1-\sum_{j=1}^{J}\left(J+p(W)^{-1}\min_{k\in\pi(j)}2^{-\iota_{X;Y}(X;Y_{k})}\right)^{-1}\biggr)\biggr]$
$\stackrel{{\scriptstyle\text{(b)}}}{{=}}$ $1-J\operatorname{\mathbb{E}}\left[\left(J+\rho(W)^{-1}\min_{1\leq k\leq L}2^{-\iota_{X;Y}(X,Y_{k})}\right)^{-1}\right]$ (81)
@@
where $(W,X,Y_{k})\sim P_{W}\times P_{X}P_{Y\mid X}$ for each $k$, (a) is by lemma 3 and (b) uses the fact that the $Y_{k}$’s are identically distributed given $X$ and the identity $\operatorname{\mathbb{E}}[\operatorname{\mathds{1}}\{\rho(W)>0\}]=\operatorname{Pr}[\rho>0]$. ∎

### v2

@@ II Problem Setup
We consider a one-shot setting with one encoder and $K$ decoders, where each decoder is connected to the encoder through a private and independent communication channel which maps an input alphabet $\mathcal{X}$ to an output alphabet $\mathcal{Y}$. Furthermore, we restrict our attention to the case where all the channels are identically modeled by a conditional distribution $P_{Y\mid X}$. Under this assumption, the encoder, defined by the mapping $f:\mathcal{W}\to\mathcal{X}$, observes the source symbol $W\in\mathcal{W}$ from $P_{W}$ and broadcasts a message $X\in\mathcal{X}$ over each channel. Subsequently, decoder $k$ views the sample $Y_{k}\in\mathcal{Y}$ at the output of its channel and generates the reconstruction $\hat{Z}_{k}\in\mathcal{Z}$ using a mapping $g:\mathcal{Y}\to\mathcal{Z}$. The probability of excess distortion for an individual decoder under the distortion measure $d:\mathcal{W}\times\mathcal{Z}\to\mathbb{R}_{\geq 0}$ is $P_{ek}=\operatorname{Pr}[d(W,\hat{Z}_{k})>D]$. As we require at least one decoder to recover the source within the specified distortion, the ensemble error probability is then $P_{e}=\operatorname{Pr}[\cap_{k=1}^{K}\{d(W,\hat{Z}_{k})>D\}]$.
We further consider a setting where each decoder has access to a conditionally independent side information realization $T_{k}\in\mathcal{T}$ from $P_{T\mid W}$, as motivated in section I. Note that for simplicity we only consider the case where the side information samples seen at the decoders are identically distributed. As opposed to the basic setting, decoder $k$ now generates $\hat{Z}_{k}$ using a mapping $g:\mathcal{Y}\times\mathcal{T}\to\mathcal{Z}$ that can exploit the side information as well as the channel observation. This constitutes a variant of the Wyner-Ziv problem [12] and consequently requires a somewhat different analysis. A block diagram of the full system model with side information is illustrated in fig. 1.
In what follows, we assume that all alphabets, e.g. $\mathcal{W}$, $\mathcal{X}$ and $\mathcal{Y}$, are standard Borel spaces endowed with their respective Borel $\sigma$-algebras and assume a common $\sigma$-finite reference measure $\theta$ on each space. For a pair of measures $\mu$, $\nu$ on $\mathcal{U}$ such that $\nu$ is absolutely continuous with respect to $\mu$, denoted as $\nu\ll\mu$, the Radon-Nikodym derivative is $d\nu/d\mu:\mathcal{U}\to\mathbb{R}_{\geq 0}$. To simplify the exposition, we assume that all probability measures to be discussed are absolutely continuous with respect to the reference measure $\theta$ on the appropriate space and work in terms of the density function $p(u)=(dP/d\theta)(u)$ for any measure $P$. Furthermore, we write $\lambda$ for the Lebesgue measure on $\mathbb{R}$ and $\lambda_{S}$ for its restriction to some $S\subseteq\mathbb{R}$. Logarithms are base-$2$ throughout; the information density for a pair of random variables $(X,Y)\sim P_{X,Y}$ is defined as
@@
[width=0.98]diagrams/system
Figure 1: System model for broadcast JSCC with side information. The source $W$ is encoded to $X$. Decoder $k$ uses the channel output $Y_{k}$ and side information $T_{k}$ to produce its reconstruction $\hat{Z}_{k}$.
@@ III-E Hybrid Strategy
$P_{e}\leq\operatorname{\mathbb{E}}\Bigl[\Bigl(1+JP_{Z}(\mathcal{B}_{D}(W))\max_{1\leq k\leq L}2^{\iota_{X;Y}(X;Y_{k})}\Bigr)^{-1}\Bigr].$ (11)
@@ IV List Poisson Matching Lemma
Let $\Pi=\{U_{i},T_{i}\}_{i\in\mathbb{N}}$ be a marked Poisson point process on $\mathcal{U}\times\mathbb{R}_{\geq 0}$ with intensity measure $\mu\times\lambda_{\mathbb{R}_{\geq 0}}$. Further associate with each point of the process an independent discrete random variable $Z_{i}$ uniformly distributed on $\{1,\ldots,K\}$ and let $P$, $Q\ll\mu$ be two probability measures on $\mathcal{U}$. Select the indices
@@
and set $U_{p}=U_{K_{p}}$ and $Z_{p}=Z_{K_{p}}$. Given $U_{p}$, let $Y_{1:K}\sim P_{Y\mid X,U}(\cdot\mid X,U_{p})$ iid. We have $(X,U_{p},Y_{k})\sim P_{X,U,Y}$ for each $k$ and note that the Markov chain condition $Y_{1:K}\leftrightarrow(X,U_{p})\leftrightarrow\{U_{i},T_{i}\}_{i\in\mathbb{N}}$ holds by construction. Next, select
@@
$\leq$ $1-\sum_{k=1}^{K}\biggl(K+\frac{p_{U\mid X}(U_{p}\mid X)}{q_{U\mid Y}(U_{p}\mid Y_{k})}\biggr)^{-1}.$ (17)
@@ V Proof of Theorem 1
First, we generate the full codebook independently from the source $W$ as a marked Poisson point process $\Pi=\{(X_{i},Z_{i}),\tau_{i}\}_{i\in\mathbb{N}}$ on $\mathcal{X}\times\mathcal{Z}\times\mathbb{R}_{\geq 0}$ with intensity measure $P_{X}\times P_{Z}\times\lambda_{\mathbb{R}_{\geq 0}}$, and associate with each point in the codebook a uniform label $C_{i}\in\{1,\ldots K\}$ independent of $\Pi$. Following the construction in [8, Th. 4], we define a modified conditional distribution $P_{\tilde{Z}\mid W}$ to take into account the intersection with the distortion ball $\mathcal{B}_{D}(w)$. The conditional density function is
@@
$\negmedspace{}\times\biggl(1-\sum_{k=1}^{K}\biggl(K+\frac{p_{X}(X)p_{\tilde{Z}\mid W}(\tilde{Z}\mid W)}{p_{X\mid Y}(X\mid Y_{k})p_{Z}(\tilde{Z})}\biggr)^{-1}\biggr)\biggr]$
@@
$\negmedspace{}\times\biggl(1-\sum_{k=1}^{K}\bigl(K+\rho(W)^{-1}2^{-\iota_{X;Y}(X;Y_{k})}\bigr)^{-1}\biggr)\biggr]$
@@
where $(W,X,Y)\sim P_{W}\times P_{X}P_{Y\mid X}$, (a) is by lemma 2 and (b) uses the fact that the $Y_{k}$’s are identically distributed given $X$ and the identity $\operatorname{\mathbb{E}}[\operatorname{\mathds{1}}\{\rho(W)>0\}]=\operatorname{Pr}[\rho(W)>0]$. ∎
@@ VI Proof of Theorem 2
We now establish the bound with side information in theorem 2. Since the side information is not visible to the encoder, it cannot restrict the codeword to the permissible distortion ball as in the proof of theorem 1. Let the codebook be a marked Poisson point process $\Pi=\{(X_{i},U_{i}),\tau_{i}\}_{i\in\mathbb{N}}$ on $\mathcal{X}\times\mathcal{U}\times\mathbb{R}_{\geq 0}$ with intensity measure $P_{X}\times P_{U}\times\lambda_{\mathbb{R}_{\geq 0}}$ independent of $W$, where $P_{U}$ is the marginal of $P_{W}P_{U\mid W}$. We associate an independent uniform label $C_{i}\in\{1,\ldots K\}$ with each point. The encoder observes $W$ and selects
@@ VIII Conclusion
We derived one-shot achievability results for a broadcast JSCC problem in a distributed setting where success is determined by the most favorable outcome among $K$ decoders. We identified two competing ways in which having multiple decoders can reduce the error probability, namely channel diversity, achieved by sharing a codebook and taking multiple independent looks at the channel, and codebook diversity, which requires forcing the decoders’ codebooks to be disjoint. During our analysis, we developed the list PML as an extension of the PML in order to obtain bounds on the error probability. Moreover, we derived a hybrid coding scheme that optimally balances channel and codebook diversity; numerical results demonstrated that it improves on approaches that rely on one form of diversity alone.
@@ Appendix A Proof of Lemma 1
Thus, an upper bound for $\operatorname{Pr}[U_{p}\neq U_{q1}\mid U_{p}=u,Z_{p}=1]$ is the probability of $\Pi_{1}^{\prime}$ seeing at least one arrival in the region
@@
$\leq$ $\frac{t}{K}\int_{\mathcal{U}}\frac{p(u)q(v)}{q(u)}\mu(dv)=\frac{tp(u)}{Kq(u)}.$ (48)
@@
$\leq$ $\operatorname{Pr}[N>0\mid U_{p}=u,Z_{p}=1]\leq 1-\biggl(1+\frac{p(u)}{Kq(u)}\biggr)^{-1}$ (52)
@@ Appendix B Proof of Lemma 2
and set $U_{p}=U_{K_{p}}$, $Z_{p}=Z_{K_{p}}$. Given $U_{p}$, let $Y_{1:K}\sim P_{Y\mid X,U}(\cdot\mid X,U_{p})$ iid. We have $(X,U_{p},Y_{k})\sim P_{X,U,Y}$ for each $k$ and note that the Markov chain condition $Y_{1:K}\leftrightarrow(X,U_{p})\leftrightarrow\{U_{i},T_{i}\}_{i\in\mathbb{N}}$ holds by construction. Next, select
@@
In what follows, condition on $\tilde{T}_{p}=t$, $U_{p}=u$, $Z_{p}=1$ and $Y_{1:K}$. Mirroring the arguments used to prove lemma 1, $\tilde{T}_{i}\geq\tilde{T}_{p}=t$ for each $i\neq K_{p}$ and therefore, by excluding the first arrival at time $t$, $\Pi^{\prime}=\{U_{i},\tilde{T}_{i}\}_{i\neq K_{p}}$ becomes a marked Poisson point process with intensity measure $P_{U\mid X}(\cdot\mid X)\times\lambda_{[t,\infty)}$ independent of $U_{p}$ and each $Z_{i}$. By the thinning property, $\Pi_{1}^{\prime}=\{U_{i},\tilde{T}_{i}\}_{i\neq K_{p},Z_{i}=1}$ is then a marked Poisson point process with intensity measure $(P_{U\mid X}(\cdot\mid X)\times\lambda_{[t,\infty)})/J$. By the Markov chain condition noted earlier, both $\Pi^{\prime}$ and $\Pi_{1}^{\prime}$ are also conditionally independent of $Y_{1:K}$ given $X$ and $U_{p}$. As in the unconditional case, assume $p_{U\mid X}(u\mid X)>0$. First consider that $q_{U\mid Y}(u\mid Y_{k})>0$ for at least one $k\in\pi(1)$. Selecting the same indices implies the selected values are the same, and consequently
@@
$\mathcal{B}=\left\{(v,s):s<\min_{1\leq k\leq L}\frac{tp_{U\mid X}(u\mid X)q_{U\mid Y}(v\mid Y_{k})}{q_{U\mid Y}(u\mid Y_{k})p_{U\mid X}(v\mid X)}\right\}.$ (64)
@@
$=$ $\frac{t}{J}\int_{\mathcal{U}}\min_{1\leq k\leq L}\max\left\{0,\frac{p_{U\mid X}(u\mid X)}{q_{U\mid Y}(u\mid Y_{k})}-\frac{p_{U\mid X}(v\mid X)}{q_{U\mid Y}(v\mid Y_{k})}\right\}$
@@
$\geq$ $\exp\left(-\frac{t}{J}\min_{1\leq k\leq L}\frac{p_{U\mid X}(u\mid X)}{q_{U\mid Y}(u\mid Y_{k})}\right).$ (73)
@@ Appendix C Error Probability of the Hybrid Scheme
In this section, we provide a proof of (11) via lemma 3. By setting $J=1$, this also serves to show the bound in (9) for the baseline scheme’s performance. Following the approach used in the proof of theorem 1, let the codebook be a marked Poisson point process $\Pi=\{(X_{i},Z_{i}),\tau_{i}\}_{i\in\mathbb{N}}$ on $\mathcal{X}\times\mathcal{Z}\times\mathbb{R}_{\geq 0}$ with intensity measure $P_{X}\times P_{Z}\times\lambda_{\mathbb{R}_{\geq 0}}$ independent of the source $W$. Associate with each point a label $C_{i}$ uniform on $\{1,\ldots,J\}$ independent of $\Pi$. The encoder uses the modified conditional distribution $P_{\tilde{Z}\mid W}$ as in (18) and selects its index $K_{p}$ using (19). Meanwhile, decoder $k$ selects $K_{qk}$ using (10). As before, let $X=X_{K_{p}}$, $\tilde{Z}=Z_{K_{p}}$, $\hat{Z}_{k}=Z_{K_{qk}}$ and $\hat{X}_{k}=X_{K_{qk}}$. The channel outputs observed by the decoders follow $Y_{1:K}\sim P_{Y\mid X}(\cdot\mid X)$ iid. We have $(W,\tilde{Z},X,Y_{k})\sim P_{W}P_{\tilde{Z}\mid W}\times P_{X}P_{Y\mid X}$ for each $k$ by lemma 3. Additionally, define $\pi(j)=\{(j-1)L+1,\ldots,jL\}$ for $1\leq j\leq J$. Then,
@@
$\negmedspace{}\times\biggl(1-\sum_{j=1}^{J}\Bigl(J+\rho(W)^{-1}\min_{k\in\pi(j)}2^{-\iota_{X;Y}(X;Y_{k})}\Bigr)^{-1}\biggr)\biggr]$
$\stackrel{{\scriptstyle\text{(b)}}}{{=}}$ $1-J\operatorname{\mathbb{E}}\Bigl[\Bigl(J+\rho(W)^{-1}\min_{1\leq k\leq L}2^{-\iota_{X;Y}(X,Y_{k})}\Bigr)^{-1}\Bigr]$ (81)
@@
where $(W,X,Y_{k})\sim P_{W}\times P_{X}P_{Y\mid X}$ for each $k$, (a) is by lemma 3 and (b) uses the fact that the $Y_{k}$’s are identically distributed given $X$ and the identity $\operatorname{\mathbb{E}}[\operatorname{\mathds{1}}\{\rho(W)>0\}]=\operatorname{Pr}[\rho(W)>0]$. ∎